If 'r' be the ratio of the roots of the equation ax^2 + bx + c = 0 , show that 
(r + 1)^2/r = b^2/ac

Grade:11

1 Answers

Arun
5 years ago
i) Let the two roots of the given quadratic equation be α & β; 
then as given α/β = r; ==> α = β*r ---------- (1) 

ii) By properties of roots of quadratic equation, 
Sum of roots: α + β = -b/a; 
Substituting from (1), α = β*r, β(1 + r) = -b/a; 
Squaring, β²(1 + r)² = b²/a² --------- (2) 

and Product of roots, α*β = c/a 
Again substituting from (1) for α = β*r, β²*r = c/a ----- (3) 

iii) Dividing (2) by (3): (1 + r)²/r = b²/ac 

Cross multiplying, {(1 + r)²}*(ac) = (b²)*(r)